a plane can fly 1120 km in 80 min with the wind. Flying against the same wind, the plane travels the same distance in 84 min. Find the speed of the wind and the speed of the plane in still air.
Found 2 solutions by TimothyLamb, stanbon:
Answer by TimothyLamb(4379)

---
x + y = 1120/80 km/min
x - y = 1120/84 km/min
---
put the system of linear equations into standard form
---
x + y = 1120/80 km/min * (60 min/hour) = 840 km/hr
x - y = 1120/84 km/min * (60 min/hour) = 800 km/hr
---
x + y = 840
x - y = 800
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x= airplane speed = 820 km/hr
y= wind speed = 20 km/hr
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php

Answer by stanbon(75887)

a plane can fly 1120 km in 80 min with the wind. Flying against the same wind, the plane travels the same distance in 84 min. Find the speed of the wind and the speed of the plane in still air.
---------------------
rate with wind:: 1120 km/(4/3)hr = 840 km/hr
-------
rate against with: 1120/(84/60) = 800 km/hr
---------------------------------
Equation:
With wind:: p + w = 840
Against wind:: p - w + 800
-----
2p = 1640
plane speed = 820 km/hr
-----
Solve for "w":
p + w = 840
820 + w = 840
w = 20 km/hr
------------------
Cheers,
Stan H.